STAFF RESOURCES

Minutes of March 9, 2005 meeting

Present:
Andy Martinez, Megan Ingegno, Mary Brodman, Irina Loutchkina, Robert Krack, Matt Spaventa, Edward Suarez, Penny Weniger, Jeff Teichmann, Eugene McElroy, Martin Katz, Melanie Robinson, Krista Dandurand, Nicole Warren, Paul Young, Stacey DeMatteo, Roselyn Riley, Shirley Peck
Absent:
Brian Stubbs, Melody Tomaszewicz

Agenda

INFORMATION SHARING

Andy/NBL- Megan Lord has been hired as the Branch Manager of SERC, and will be coming around to most units as part of her training. Meghan’s email is mlord@rci.rutgers.edu.

Paul/Gov Docs- A mini-shift is being planned to accommodate the FCC Record and the project is expected to take several months to complete.

Megan/Physics- Has completed phase I of the Serials Inventory Project and preparations are being made for a proposed weeding project and shift of the bound periodicals collection.

Matt/Douglass- Preparing for a shift of the East Room Stacks Collection during Spring Break. Also working on withdrawing approximately 100 items from the Stacks Collection in the East Room.

Eugene/SMLR- Has been in contact with a staff person from the Ceramics Department to acquire a 3M-security gate at a discounted price, he anticipates purchase and installation should be soon.

Eddie/LSM- Has completed the design of building floor maps, to help patrons navigate the library. Working on the packing and shipping of the final items left in the LSM Special Collections Room. Should be completed by the beginning of April/2005.

Melanie/Math- Progress continues on the Serials Inventory project (part time employee has been trained to assist on the project). Collection space is being assessed with an eye towards what collections are growing at a faster rate. Approximately 500 bound periodicals will be transferred to the Annex after work is completed on inventory and line adding. The transfer should occur towards the end of May/2005.

Jeff/Alex- Pricilla Lee is now responsible for RDS duties, since the departure of Will Torres. The assistant to the student coordinator position is now vacant and has been posted.

Stacy/Annex- Has received 186 Protein Database boxes to be stored as part of a National Science Foundation and Chemistry Department project. Continues to receive about 400 items per week in transfers from the LSM Weeding Project.

Irina/Alex- Reserve processing has slowed, so she has been helping Shirley with In- Processing. Also continues to repair damaged reserve materials as needed.

Mary/Alex- The collection of Rona Goffen (deceased RU Art History Professor) was donated to the libraries. Items have been unboxed and shelved and now await review by selectors. Some items will remain at Alex and some may go to the Art Library. Weeding of the Pane collection has been completed; some items will go the Africa and some to LSM Special Collections as un- cataloged display items.

Rob/Alex- As part of water emergency preparedness, rolls of plastic sheeting from water emergency kits were cut into three different sizes to fit over ranges. A mini shift of 2A is in progress and in the near future mini shifts are planned for 2B, 3 and the basement to accommodate the increase of new additions to the collection.

Shirley/Alex- Working on searching claims returned lists and has had success finding items. There’s been an increase in In-Process books, many of which will fill the New Books shelf near the lobby entrance.

STANDARD RANGE GUIDE TEMPLATES/INSTRUCTIONS

The final template and instructions have been approved and are now available on the NBL Access Services page under the Collection Management documentation heading. (NBL Access Services page/Functional Groups/Collection Management-Documentation/Range Guide Information)

The instructions are posted as a PDF document while the two separate templates (LC Class Number and Shelved by Title) are posted as Microsoft Word documents. Instructions are posted with the template headings on how to download the templates to your computer.

The group was asked to review the documents and report any problems accessing or using the template to Andy. Several shifting projects are in progress or planned for the near future so the templates will be in use shortly.

LC CLASSIFICATION DOCUMENTATION

Rob and Eddie are working on revising the documentation currently available on the NBL page (NBL Access Services page/Functional Groups/Collection Management-Documentation/Collection Management-Library of Congress Classification System)

The consensus is that it needs to be condensed and updated to included a different call number test (maybe separate from document)

Rob will also create documentation of the successful procedures used at Alex for training students on the systems, and distribute it to group members

Rob and Eddie’s revisions will be reviewed at the next meeting

COLLECTION FLOOR MAPS

Floor maps of the different units were circulated for review and will be revised to also include a focus on building security.

It’s been proposed that a floor map handout will be posted and distributed in the units for patron reference on site, and a more in-depth electronic resource will also be made on the web

A subgroup will be formed to work on this project, comprised of members from the Collection Management group and the Building Security and Maintenance group. The first meeting of the sub-group will be announced at a later date. If you wish to volunteer to work on this project, please let Andy or Penny know.

EAST ROOM SHIFT AT DOUGLASS LIBRARY

During the week of March 15, 2005 (Spring Break), the East Room and part of the Tapestry Room at Douglass will be shifted to maximize the available space and also empty some shelving to be moved to another unit.

Matt Spaventa, along with 3 work-study students and volunteers from the group, will work on the shift

Once the shift is completed, the Music Library will be able to use some of the shelving emptied in the Tapestry Room for their library. 10 double-sided sections of shelving will be moved to the Music Library.

Mary made workflow suggestions to help the project progress more efficiently

A discussion about the cleaning of the shelving arose, in regard to which cleaning products would be safest to use without causing damage to the books. After consultation with Collection Services, it was decided that warm water would be used with rags to wipe away the dust. Paper towels will be used in case excess water needs to be wiped away. The approximate time that it takes from start to finish to shift each range with the incorporation of shelf cleaning will be noted during the project. These times will let us know about the future feasibility of incorporating shelf cleaning into large shifts.

TEMPERATURE/HUMIDITY READERS IN NBL

LCD monitors and data loggers that take humidity and temperature level readings will be installed in all of the NBL unit. Each unit will have some combination of the two. The equipment will allow for constant monitoring of humidity and temperature levels and will allow us to get levels over time.

Once all of the equipment is installed and running, Ian Bogus will visit the group to discuss the project.

NEXT MEETING

Wednesday April 13, 2005
2:30 – 4:30 pm
Location to be announced



 
URL: http://www.libraries.rutgers.edu/rul/staff/groups/collection_mgt_group/minutes/collection_mgt_group_05_03_09.shtml
Website Feedback  |  Privacy Policy

© Copyright 1997-2013, Rutgers University Libraries   (Further Copyright Information)